I need help with Literal Equations, my textbook doesn't explain it very good. One of the equations is (x/j-k)-j=k and I'm solving for "x". I put the parentheses to show that it was a fraction, they aren't there in my book. If you could help me I would greatly appreciate it.
Answer by stanbon(75887)

(x/(j-k))-j=k and I'm solving for "x"
-----
x/(j-k) = j+k
---
Multiply both sides by (j-k) to get:
x = (j-k)(j+k)
---
x = j^2-k^2
=================
Cheers,
Stan H.